OpenStudy (anonymous):

A helium-4 atom falls from a height h into a tank of gaseous helium atoms at 300 K. What is h in meters if the temperature of the helium atoms remains at 300 K after the falling atom enters the tank? Details and assumptions The acceleration of gravity is −9.8 m/s2. As the atom falls it is in vacuum. The atom starts from rest. please help

OpenStudy (anonymous):

how should i approach this problem

OpenStudy (chmvijay):

consider helium to be like stone OK

OpenStudy (chmvijay):

since its falling on the earth g=*9.8

OpenStudy (chmvijay):

4 helium atoms are there or only one Helium atom is there

OpenStudy (anonymous):

helium 4 means the mass number 2 proton and 2 electrons

OpenStudy (chmvijay):

ok then you got the mass of the stone too

OpenStudy (anonymous):

THis it was initially at rest it was at 0k,then mgh=mcT

OpenStudy (anonymous):

mass isn't given it should be done based on the assumption above

OpenStudy (chmvijay):

yaaa mass u take as 4 gram ok

OpenStudy (chmvijay):

what is McT

OpenStudy (anonymous):

it is the formula fo heat transfer,m is mass,c is specific heat capacity and t is the temperature,the mgh=mct ,gh=ct, i have no idea how to get the specific heat capacity,

OpenStudy (chmvijay):

they have asked to find the height dude why are you relating to specific heat capacity

OpenStudy (anonymous):

because there is a temperature change from 0k to 300k.the mass isn't given

OpenStudy (anonymous):

and mechanical work causes a change in temperature

OpenStudy (chmvijay):

they have given constant temperature of 300 K LOL:)

OpenStudy (anonymous):

the rms velocity of an atom is given by \[v=\sqrt{3RT/M}\] R=gas constant T=temperature M=molar mass For helium M = 4 grams and it is given that it enters with such velocity that its temperature is 300K. We can thus know the velocity. Hence the kinetic energy which was initially stored as potential energy. Apply conservation of mechanical energy and find h.

OpenStudy (chmvijay):

mgh =mvh ????or what

OpenStudy (anonymous):

mgh = mv^2/2 where v is the velocity we found from above formula

OpenStudy (chmvijay):

ok ok i was getting trouble to correlate the right side term any way thank you

OpenStudy (anonymous):

a single atom can not weigh 4 grams,

OpenStudy (chmvijay):

h=v^2/2g

OpenStudy (anonymous):

no.if an atom is at rest,it means it is at 0k,absolute zero

OpenStudy (chmvijay):

LOL:( no its not OK K is unit of temperature how can you relate this too rest

OpenStudy (anonymous):

is 11.47 correct ?

OpenStudy (anonymous):

because absolute zero is the temperature at which there is no vibration or movement,hence no kinetic energy

OpenStudy (chmvijay):

in this case m will cancel and not needed LOL:)

OpenStudy (anonymous):

how do u get 11.47

OpenStudy (anonymous):

so what's the answer ?

OpenStudy (chmvijay):

V2/2g

OpenStudy (anonymous):

i equated the K.E. and the P.E. taking m = 4 grams ! is it correct ?

OpenStudy (anonymous):

no,4 grams is way to big for a single atom

OpenStudy (anonymous):

so now what is the correct answer ?

OpenStudy (chmvijay):

just solve the square root of 3RT/M putting T =300K ,M=4 , R=8.314

OpenStudy (anonymous):

that will give the rms velocity not height !

OpenStudy (anonymous):

Q=mct,where m is the mass,c is heat capacity and t is change in temperature. where q = mgh thus mgh=mct,which m cancels out,and it simplify it to h 300c/9.8, c=3/2r for monoatomic gases

OpenStudy (chmvijay):

v=43.25 next find the h

OpenStudy (chmvijay):

h=V2/2g

OpenStudy (chmvijay):

95.40 meters

OpenStudy (anonymous):

if r= 3/2r where r is gas constant,so h = 300*3/9.8*2r

OpenStudy (chmvijay):